Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120816

Observers : C. Inserra (QUB) Flora Cellier-Holzem (LPHNE) 
Data analysts on duty : M. Fraser, D. Wright, A. Pastorello

Time shared with Benetti Large Program (BLP) also over the second three nights rather than the breakdown of firewalled nights - this was due to RA pressue ranges of targets and having to work around the weather. Data reduced, as much as possible with the PESSTO pipeline

Total time spent BLP on 20120816 : 0h 
Total time overall in Aug 2012 : 16h (out of total of about 31hr clear) 
Total time lost on 20120816: 10h 

All times CST and approximate
Bad weather since yesterday: storm, clouds and strong wind!!

21.30 Clouds everywhere, still closed
00.30 Still closed
03.30 Sky with some stars. Still closed due to a combination of dark clouds and fog.
05:40 Dome still close.
06.00 Twilight incoming, we are leaving the building. Stop here!!
Comments